OpenStudy (idealist10):

Solve y'=(-6x+y-3)/(2x-y-1) implicitly.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Equations like these tend to simplify nicely when you try a substitution...

OpenStudy (idealist10):

What substitution?

OpenStudy (anonymous):

Currently trying to figure that out. I've tried a few already but they don't seem to help.

OpenStudy (anonymous):

I'll be back in a while... maybe some idea will crop up.

OpenStudy (idealist10):

Any ideas?

OpenStudy (anonymous):

Actually, yes. I found a neat resource online for dealing with equations like this. First we let \(X=x-x_0\) and \(Y=y-y_0\). Then we have \(y'=Y'\). \[\begin{align*}Y'&=\frac{-6(X+x_0)+(Y+y_0)-3}{2(X+x_0)-(Y+y_0)-1}\\\\ Y'&=\frac{-6X+Y+(-6x_0+y_0-3)}{2X-Y+(2x_0-y_0-1)}\end{align*}\] What we want to do is find constants \(x_0\) and \(y_0\) such that the constant terms in the numerator and denominator are both zero; the resulting equation will be easier to deal with. \[\begin{cases}-6x_0+y_0-3=0\\ 2x_0-y_0-1=0\end{cases}~~\implies~~x_0=-1,~y_0=-3\] You then have \[\begin{align*}Y'&=\frac{-6X+Y}{2X-Y}\\\\ Y'&=\frac{-6+\dfrac{Y}{X}}{2-\dfrac{Y}{X}}\end{align*}\] Substitute \(Y=uX\), so \(Y'=Xu'+u\). \[\begin{align*}Xu'+u&=\frac{-6+u}{2-u}\\\\ Xu'&=\frac{-6+u}{2-u}-u\\\\ Xu'&=\frac{-6+u-u(2-u)}{2-u}\\\\ Xu'&=\frac{-6-u+u^2}{2-u}\\\\ \frac{2-u}{u^2-u-6}~du&=\frac{dX}{X}\end{align*}\]

OpenStudy (idealist10):

u=x+h v=y+k dy/dx=dv/du dv/du=(-6(u-h)+v-k-3)/(2(u-h)-(v-k)-1) dv/du=(-6u+v+(6h-k-3))/(2u-v+(-2h+k-1)) 6h-k-3=0 -2h+k-1=0 so h=1, k=3 dv/du=(-6u+v)/(2u-v) dv/du=(-6+v/u)/(2-v/u) v=uz dv/du=u*dz/du+z u*dz/du+z=(-6+z)/(2-z) u*dz/du=(-6+z)/(2-z)-z u*dz/du=(-6+z)/(2-z)-(z(2-z))/(2-z) =(-6+z-2z+z^2)/(2-z) =(-6-z+z^2)/(2-z) =(z^2-z-6)/(2-z) =((z-3)(z+2))/(2-z) (2-z)/((z-3)(z+2)) dz=du/u am I right so far?

OpenStudy (anonymous):

Yes

OpenStudy (idealist10):

Then can you tell me the right answer? Because the answer I got is (y-3x)(y+2x+5)^4=C.

OpenStudy (anonymous):

Here's what I'm getting:\[\begin{align*}\frac{2-u}{u^2-u-6}~du&=\frac{dX}{X}\\\\ \int\frac{2-u}{(u+2)(u-3)}~du&=\int\frac{dX}{X}\\\\ \int\frac{2-u}{(u+2)(u-3)}~du&=\ln|X|+C\\\\ -\frac{1}{5}\int\frac{du}{u-3}-\frac{4}{5}\int\frac{du}{u+2}&=\ln|X|+C\\\\ -\frac{1}{5}\left(\ln|u-3|+4\ln|u+2|\right)&=\ln|X|+C\\\\ \left((u-3)(u+2)^2\right)^{-1/5}&=CX\\\\ (u-3)(u+2)^2&=CX^{-5}\\\\ (u-3)(u+2)^2&=CX^{-5}\\\\ \left(\frac{Y}{X}-3\right)\left(\frac{Y}{X}+2\right)^2&=CX^{-5}\\\\ \left(\frac{y+3}{x+1}-3\right)\left(\frac{y+3}{x+1}+2\right)^2&=C(x+1)^{-5}\\\\ \left(\frac{y-3x}{x+1}\right)\left(\frac{y+2x+5}{x+1}\right)^2&=C(x+1)^{-5}\\\\ \left(y-3x\right)\left(y+2x+5\right)^2&=C(x+1)^{-2} \end{align*}\]

OpenStudy (idealist10):

I got (z-3)(z+2)^4=C/u^5

OpenStudy (anonymous):

Let me see if I get a different result on paper... I tend to make slight mistakes when I only use LaTeX.

OpenStudy (idealist10):

I think you got it wrong in the middle. You have (u+2)^2 but I have ^4 since it's 4*ln abs(z+2).

OpenStudy (anonymous):

Ah there it is. Right, so it should be \[(u-3)(u+2)^4=CX^{-5}\] where you use \(z\) in place of my \(u\) and \(u\) in place of my \(X\), right?

OpenStudy (idealist10):

Yes.

OpenStudy (anonymous):

\[\left(y-3x\right)\left(y+2x+5\right)^4=C\]

OpenStudy (idealist10):

So I got it right?

OpenStudy (idealist10):